ILovePuppiesLol:

describe an infinite geometric series with a beginning value of 2 that converges to 10. what are the first 4 terms of the series

ILovePuppiesLol:

@Shadow

bm717:

a{1} = first term of series ∞ Infinite Sum =∑ a{1} • r^(n–1) = a{1} ⁄ (1–r)...for any geometric series n=1 Infinite Sum for this problem = 10 = a{1} ⁄ (1–r)...a{1} = 2 (given) 10 = 2 ⁄ (1–r) r = 0.8...common ratio ∞ Infinite Sum =∑ 2 • (0.8)^(n–1) ◀◀ (answer) n=1 a{n}=2 • (0.8)^(n–1) a{1}=2 • (0.8)^(1–1) = 2 a{2}=2 • (0.8)^(2–1) = 1.6 a{3}=2 • (0.8)^(3–1) = 1.28 a{4}=2 • (0.8)^(4–1) = 1.024 a{5}=2 • (0.8)^(5–1) = 0.8192 Infinite Sum =2 + 1.6 + 1.28 + 1.024 + 0.8192 + . . . + 2•(0.8)^(n–1) ◀◀ (answer)

ILovePuppiesLol:

did u just copy and paste

bm717:

Mhm (:

ILovePuppiesLol:

Cheese

ILovePuppiesLol:

@Shadow

bm717:

dw it is right